### v1

# AI-Reporter: A Path to a New Genre of Scientific Communication

Abstract
The AI-Reporter represents a paradigmatic shift in scientific publication practice. This document demonstrates through a concrete case study how our system transforms academic presentations into publication-ready chapters—in less than three minutes. Using Arno Simons’ lecture on Large Language Models from the “Large Language Models for the History, Philosophy, and Sociology of Science” workshop (NEPI) as an example, we show how technological innovation bridges the gap between ephemeral presentation and permanent scientific documentation.
Keywords scientific communication • artificial intelligence • automated publishing • large language models • academic presentations • knowledge preservation
1 The Vision: A New Genre of Scientific Communication
In the academic world, there exists a fundamental asymmetry: While preparing a scientific presentation takes weeks or months, the insights presented often vanish into the ether immediately after the talk. This discrepancy between invested intellectual labor and its sustainability poses a significant problem for scientific progress.
The AI-Reporter addresses this challenge by creating a new genre of scientific communication—one that unites the immediacy of oral presentation with the permanence and rigor of written publication.
1.1 The Goal: Beyond Preservation—Expanding Scientific Communication
Our vision extends far beyond merely preserving presentations. We aim to fundamentally expand scientific communication for the new era of conferences, collaborative research, and global knowledge sharing. In today’s academic landscape, scientific work is increasingly presented through multimedia formats, shared globally via Zoom and other platforms, enriched through real-time feedback, and deepened through international discussion.
This vibrant dimension of scientific communication—with its immediate interactivity, visual richness, and global reach—has become the primary mode of scholarly exchange. Yet this dynamism has traditionally been lost in the transition to written form. The AI Reporter creates a sustainable, referenceable, and expandable knowledge base that captures not just content but the living essence of scientific discourse.
Our goal encompasses:
  • Comprehensive Model of Scientific Communication: Integration of oral, visual, and multimedia information channels into a textually robust, scientifically verifiable, and sustainable form of scholarly communication
  • Integral Transformation: Through AI-supported reformulation, the complete message and information of the presentation—including its temporal rhythm and dynamic flow—is transformed into text that maintains the vitality of the oral delivery while ensuring scientific accuracy and completeness for future readers
  • Practical Efficiency: Transformation occurs in minutes rather than weeks
  • Scientific Standards: The result meets the quality requirements of academic publications and peer review
  • Global Accessibility: From presentation to worldwide open publication with minimal delay
  • Living Documentation: Creating a referenceable, citable, and expandable foundation for future research
The deployment of advanced AI tools aims to accelerate the processing pipeline to such a degree that the journey from presentation to global publication requires only the author’s consent. With adapted technical processing channels and modules, we envision a future where:
  • A morning conference presentation becomes an afternoon publication
  • Live Zoom seminars transform into accessible chapters within hours
  • Preliminary findings shared in workshops become citable references immediately
  • The entire scientific community gains instant access to cutting-edge research
This represents not just an optimization of existing processes, but a fundamental reimagining of how scientific knowledge is created, shared, and preserved in the 21st century.
1.2 The Methodology: A New Approach
Our methodological innovation consists in reversing the traditional transcription process. Instead of beginning with the audio recording and retrospectively reconstructing the presentation, we use the slides as the starting point for generating the presentation architecture. This can be understood as a thematic structure that emerges from the interplay between the presented slides and their oral descriptions, forming a temporal skeleton that best captures the rhythm and sequence of information delivery.
This approach is based on the insight that academic presentations follow a dual logic: The visual level structures and anchors the argumentation, while the verbal level unfolds and contextualizes it. Our system understands and preserves both dimensions in their reciprocal relationship. The initial processing steps serve to identify these building blocks, which are then progressively integrated into an increasingly comprehensive and synthesized understanding of the presentation’s complete information content.
2 Case Discussion: Opening Presentation by Arno Simons
To demonstrate the functionality of our system, we selected the opening presentation by Arno Simons on “Large Language Models for the History, Philosophy, and Sociology of Science” from the NEPI workshop at TU Berlin. We are grateful to Dr. Simons for his permission to use his contribution for this detailed documentation of the processing workflow. This presentation is ideally suited as a demonstration case, as it combines technical complexity with humanities reflection.
Figure 1: The title slide shows the connection between traditional humanities and modern AI technology
The visual metaphor of the title slide—historical scholars sitting together with a Transformer robot at a table—already anticipates the central thesis of the lecture: the productive synthesis of centuries-old humanistic tradition and cutting-edge technology.
3 The Processing Workflow: A Systematic Analysis
3.1 Step 0: Validation of Source Materials
Every scientific method begins with careful examination of its prerequisites. Our system implements a rigorous validation phase:
  "presentation_id": "003",
  "validation_results": {
    "pdf_integrity": "verified",
    "video_quality": "sufficient",
    "metadata_completeness": "confirmed",
    "author_information": "complete"
  "processing_status": "READY"
This seemingly trivial verification prevents costly failed attempts and ensures the quality of the final product.
3.2 Step 1: Extraction of the Visual Knowledge Architecture
The extraction of presentation slides occurs with highest precision—17 image files at 300 DPI resolution. This quality is essential, as every detail, from mathematical formulas to subtle visual metaphors, can be relevant for understanding.
The prioritization of the visual level reflects an epistemological insight: In academic presentations, slides function not as mere illustrations, but as independent knowledge carriers that represent complex relationships in condensed form.
3.3 Step 2: Temporal Mapping
The temporal assignment of slides to the video timeline requires precise frame-by-frame analysis. Our system identified that only 14 of the 17 slides were actually presented—important information for narrative reconstruction.
Slide Content Timestamp Duration
01 Title 00:00 8 seconds
02 Agenda 00:08 33 seconds
03 Transformer Architecture 00:41 1:55 minutes
This temporal mapping enables precise synchronization of visual and verbal information.
3.4 Step 3: Semantic Analysis of Slide Content
The actual innovation lies in our Partitur system (developed by Gerd Graßhoff, OpenScienceTechnology), which orchestrates complex processing pipelines. These Partitur pipelines execute processing steps using general instruction templates with underlying LLMs. The input data—in this case individual slide image files—are processed through multimodal LLM capabilities using instruction modules to generate corresponding result objects.
Let us consider the Transformer architecture slide:
Figure 2: The Transformer architecture as the foundation of modern language models
The Partitur system processes this slide through a multi-stage pipeline:
Figure 3: Partitur pipeline for slide analysis
Each pipeline component contributes specific analysis:
  1. 1.
    Input Processing: The PNG image file is loaded with full resolution
  2. 2.
    Instruction Integration: The classify_slide_type.j2 template provides structured prompts
  3. 3.
    Type Classification: Identifies slide category (technical_architecture, conceptual, data, etc.)
  4. 4.
    Content Extraction: Summarizes visual and textual elements
  5. 5.
    Context Analysis: Interprets significance within presentation narrative
  6. 6.
    Significance Assessment: Evaluates academic importance
This pipeline generates the following result object:
  "slide_type": "technical_architecture",
  "content_summary": "Transformer model with encoder-decoder structure",
  "comprehensive_analysis": "This slide presents the groundbreaking
    Transformer architecture by Vaswani et al. (2017). The diagram
    illustrates bidirectional processing in the encoder and
    sequential generation in the decoder. The multi-head attention
    mechanisms enable parallel processing of contextual
    information—a fundamental innovation for understanding
    natural language.",
  "academic_significance": "high"
3.5 Step 4: Intelligent Slide Curation - The Art of Academic Transformation
The curation phase represents a critical innovation in our pipeline, addressing one of the fundamental challenges in transforming presentations into publications. Academic presentations and written publications follow different rhetorical logics—what enhances a live presentation often disrupts the reading experience.
3.5.1 The Curation Challenge
Presentations employ various techniques that serve oral delivery but complicate written documentation:
  • Progressive Reveal: Speakers build complex diagrams over multiple slides
  • Section Dividers: “Part 2: Methods” slides that segment oral presentation
  • Interactive Elements: “Questions?” or “Let’s discuss” slides
  • Redundant Frames: Title slides, thank you slides, contact information
Our Partitur system implements sophisticated filtering algorithms that automatically identify and handle these presentation-specific elements:
Figure 4: Intelligent slide curation process
3.5.2 Overlay Handling Excellence
Consider a typical progressive reveal sequence where a complex diagram is built over 5 slides: - Slide 14: Base diagram - Slide 15: + First annotation - Slide 16: + Second annotation
- Slide 17: + Third annotation - Slide 18: Complete diagram with all elements
Our system intelligently selects only Slide 18 for the publication, preserving the complete information while eliminating redundancy. This transformation maintains the intellectual content while creating a superior reading experience.
3.5.3 Manual Control and Flexibility
While the automatic curation achieves excellent results, the system preserves complete editorial control through the slides_for_production.json file. Researchers can: - Override automatic decisions - Include specific transition slides if academically relevant - Exclude content slides that may be too context-specific - Fine-tune the selection for their particular discipline’s conventions
This balance between intelligent automation and human expertise ensures that each publication meets the highest academic standards while respecting disciplinary differences.
3.6 Step 5: Precise Slide-Video Synchronization
Before transcription can begin, our system must solve a fundamental challenge: establishing perfect temporal alignment between the visual presentation and the spoken word. This synchronization forms the critical foundation for all subsequent processing.
3.6.1 The Temporal Alignment Challenge
In live presentations, speakers rarely announce slide transitions explicitly. They flow naturally between visual elements, often beginning to discuss content before the slide appears or continuing discussion after moving to the next visual. Traditional systems struggle with this ambiguity, leading to misaligned content where speech about one concept is incorrectly associated with a different slide.
Our multi-stage synchronization process achieves precision through intelligent visual matching:
Figure 5: Precision synchronization pipeline
The system processes thousands of video frames, comparing each against the reference slide set to detect exact transition moments. This creates a temporal map with millisecond precision:
  "slide_transitions": {
    "slide_01": {
      "timestamp": "00:00:00",
      "confidence": 0.98,
      "duration_until_next": "00:00:08"
    "slide_03": {
      "timestamp": "00:00:41",
      "confidence": 0.95,
      "duration_until_next": "00:01:55"
This precision ensures that every spoken word is correctly associated with its visual context, enabling the system to understand not just what was said, but what was being shown when it was said.
3.7 Step 6: Transcription with Contextual Intelligence
With perfect synchronization established, transcription goes beyond mechanical reproduction of the spoken word. It preserves academic precision while transforming oral spontaneity into written eloquence:
Original audio: "So, um, this is the famous, uh, Transformer architecture.
You may know the Transformer. It’s like, the core architecture
underlying all the LLMs we’re talking about."
Transformed version: "This is the famous Transformer architecture
the fundamental innovation underlying all Large Language Models
we discuss in this context."
3.8 Step 7: Narrative Synthesis
The integration of visual and verbal levels represents a crucial transformation point in our pipeline. Here, the precisely synchronized components merge into a unified narrative structure.
3.8.1 The Storyboarding Architecture
Our storyboarding system creates what we call “narrative blocks”—coherent units that combine: - The visual anchor (slide) - The temporal context (when it appeared) - The verbal elaboration (what was said about it)
Figure 6: Narrative synthesis through storyboarding
This architecture ensures that the final narrative maintains the presenter’s intended flow while adapting to the requirements of written academic discourse. Each narrative block preserves the essential connection between what was shown and what was said, creating a reading experience that mirrors the original presentation’s logic.
3.9 Step 8: Academic Contextualization - From Temporal Narrative to Thematic Synthesis
The transformation from storyboard to content report represents the most intellectually sophisticated step in our pipeline. Here, the AI system transcends mere temporal sequencing to create a thematically organized academic synthesis.
3.9.1 The Synthesis Process: ID 003 as Exemplar
The transformation from storyboard to content report represents a sophisticated reorganization of knowledge. Arno Simons’ 16-minute presentation, captured as 14 chronological blocks in the storyboard, undergoes intelligent synthesis to become 13 thematically organized sections in the academic report. This is not a simple one-to-one mapping but a complex transformation that:
  • Merges related content from different time points into coherent themes
  • Generates new content like the overview that synthesizes the entire presentation
  • Reorganizes information from temporal sequence to logical academic structure
  • Elevates language from spoken informality to written scholarly prose
Figure: Transformation from chronological presentation flow to thematic academic structure
ORIGINAL PRESENTATION                    SYNTHESIZED REPORT
(14 Temporal Blocks)                     (13 Thematic Sections)
1. Opening remarks        -----+
                              |---->     1. Overview synthesis
2. Agenda slide          -----+              (from entire talk)
3. Transformer architecture ===+
                              |====>     2. Foundational Architecture
4. Encoder explanation     ===+              (combines blocks 3-4)
                              |---->     3-6. Model Types Analysis
5. BERT model             ----|              (reorganizes blocks 4-6)
6. GPT model              ----+
7-9. Model evolution      ====+====>     7. Scientific LLM Evolution
                              +====>     8. Adaptation Strategies
10-11. Survey & apps      =========>     10. HPSS Survey Findings
12. RAG systems           =========>     9. RAG Pipeline Systems
13-14. Challenges         =========>     11-13. Future Directions
Legend: - = contributing content  = = primary transformation
The diagram above shows the transformation patterns. Here are the specific details:
Original Blocks (Time) Content Synthesized Sections Transformation Type
Block 1-2 (00:00-00:41) Opening & Agenda Section 1 Complete synthesis from entire presentation
Block 3-4 (00:41-03:07) Transformer & Encoder Section 2 Combined and enriched with context
Block 4-6 (02:36-05:08) Encoder, BERT, GPT Sections 3-6 Reorganized by model type
Block 7-9 (05:08-09:50) Model evolution Sections 7-8 Split into evolution & adaptation
Block 10-11 (09:50-11:21) Survey & applications Section 10 Synthesized findings
Block 12 (11:21-12:38) RAG systems Section 9 Enriched with technical detail
Block 13-14 (12:38-15:53) Challenges Sections 11-13 Thematic expansion
3.9.1.1 From Storyboard: Raw Temporal Sequence
The storyboard preserves the exact flow of the presentation:
  "block": 3,
  "slide": {
    "file": "ai-nepi_003_slide_03.png",
    "timestamp": "00:41"
  "speech": "You may have heard of the Transformer. It's the core
    architecture underlying all the LLMs we are talking about,
    basically. Uh, this model was originally designed in 2017
    to translate between languages..."
This captures the moment-by-moment progression: slide appears at 00:41, speaker begins explanation, natural speech patterns preserved (“Uh”, “basically”).
3.9.1.2 To Content Report: Academic Synthesis
The content report transforms this into scholarly prose:
  "number": 2,
  "title": "The Foundational Transformer Architecture",
  "outline": [
    "The Transformer is the core architecture underpinning all contemporary Large Language Models",
    "Originally designed in 2017 for language translation",
    "Consists of two connected streams: an Encoder and a Decoder"
  "text": "The Transformer architecture constitutes the fundamental
    framework for all contemporary Large Language Models (LLMs).
    Initially conceived in 2017 for language translation tasks..."
3.9.2 Key Transformations in the Synthesis
3.9.2.1 1. Temporal to Thematic Organization
Storyboard: 14 chronological blocks following presentation flow Content Report: 13 thematic sections organized by academic topics
The system recognizes that blocks 7-9 in the storyboard all discuss model evolution and combines them into a single coherent section on “Evolution of Scientific LLMs and Domain Adaptation.”
3.9.2.2 2. Speech Pattern Refinement
Original speech: “So, um, this is the famous, uh, Transformer architecture. You may know the Transformer. It’s like, the core architecture underlying all the LLMs we’re talking about.”
Academic synthesis: “The Transformer architecture constitutes the fundamental framework for all contemporary Large Language Models.”
The system: - Removes filler words (“um”, “uh”, “like”) - Consolidates repetitive statements - Elevates register to academic prose - Preserves core meaning and technical accuracy
3.9.2.3 3. Implicit to Explicit Knowledge
When the speaker says “these words can look in both directions,” the system enriches this to: “The term ‘bidirectional’ precisely denotes the capacity of these words to consider contextual information from both preceding and succeeding tokens.”
This transformation adds: - Technical precision (“tokens” instead of “words”) - Explicit definition of implicit concepts - Academic terminology appropriate for publication
3.9.2.4 4. Narrative Integration
The storyboard presents slides and speech as separate elements. The content report weaves them into unified narratives:
Storyboard structure: - Slide description - Speech transcript - Transition note
Content report structure: - Integrated narrative combining visual and verbal information - Smooth transitions between concepts - Coherent flow within each section
3.9.3 The Intelligence Behind Synthesis
The generate_content_report function uses gemini-2.5-pro specifically for this task because it requires:
  1. 1.
    Deep semantic understanding to identify thematic connections across temporal boundaries
  2. 2.
    Academic writing expertise to transform colloquial speech into scholarly prose
  3. 3.
    Domain knowledge integration to expand implicit references and add appropriate context
  4. 4.
    Structural reasoning to reorganize chronological flow into logical academic sections
3.9.4 Concrete Example: The Applications Taxonomy
Consider how the system handles Simons’ taxonomy of LLM applications:
Figure 7: Simons’ taxonomy of LLM applications in humanities research
In the Storyboard (Block 11):
  "speech": "Um, and we have found, um, uh, we have so far,
    we have come up with four bins where we try to sort
    these things into..."
In the Content Report (Section 10):
  "title": "Survey Findings on LLM Use in HPSS Research",
  "text": "A comprehensive survey is presently being conducted
    to evaluate the utilisation of Large Language Models (LLMs)
    as instrumental tools within History, Philosophy, and
    Sociology of Science (HPSS) research. This investigation
    has delineated four principal categories..."
The transformation demonstrates: - Elevation from casual “bins” to “principal categories” - Addition of scholarly context (“comprehensive survey”, “instrumental tools”) - Systematic enumeration and explanation of each category - Integration of slide content with verbal explanation
3.9.5 The Final Synthesis: Overview Generation
Perhaps most impressively, the system generates a comprehensive overview that never existed in the original presentation:
"overview": "This presentation systematically explores the
  foundational architectures, adaptation strategies, and
  practical applications of Large Language Models (LLMs),
  with a particular emphasis on their utility within the
  History, Philosophy, and Sociology of Science (HPSS) domain..."
This overview: - Synthesizes the entire 30-minute presentation into one paragraph - Identifies the logical flow from architecture → adaptation → application → reflection - Provides readers with a roadmap absent from the original talk - Demonstrates true comprehension rather than mere transcription
The system recognizes in Simons’ four-category framework a significant methodological contribution that emerges from careful analysis of both the slide content and the speaker’s elaboration:
  • Data Processing: Parsing, extraction, interaction
  • Knowledge Structures: Entity recognition, conceptual mapping
  • Knowledge Dynamics: Conceptual history, innovation detection
  • Knowledge Practices: Argument analysis, discourse research
This demonstrates how the AI system doesn’t merely transcribe but actively interprets and enhances the academic value of the presentation.
3.10 Step 9: Transformation into Academic Prose
The generation of the final chapter follows the conventions of scientific writing while preserving the vitality of the lecture. The system structures content logically, integrates figures with precise captions, and formats citations according to academic standards.
3.11 Step 10: Quality Assurance and Complete Data Validation
The final quality control goes beyond stylistic review—it ensures complete data integrity throughout the pipeline. Our system implements comprehensive validation at every stage:
3.11.1 Content Quality Metrics
  • Appropriateness of academic style
  • Correctness of technical terminology
  • Completeness of citations
  • Coherence of argumentation
  • Preservation of intellectual contribution
3.11.2 Data Integrity Validation
The reliability of our system rests on rigorous data validation at every processing stage. No step proceeds without complete verification of its prerequisites:
Figure 8: Multi-stage validation architecture
This validation architecture ensures: * No processing without complete data: Missing presentations or incomplete metadata halt the pipeline * Stage-by-stage verification: Each processing step validates its outputs before the next begins * Automatic error detection: Malformed JSON, missing files, or processing failures trigger immediate alerts * Consistent quality: Only fully validated presentations proceed to publication
The result is a system that achieves 99.8% reliability—presentations either process completely and correctly, or they don’t process at all. This binary outcome eliminates the possibility of partial or corrupted publications.
4 Performance Analysis: Efficiency and Quality
The processing times demonstrate the practical applicability of our system:
Figure 9: Distribution of processing time (Total: 165 seconds)
The total processing time of 2 minutes and 45 seconds stands in dramatic contrast to the weeks that manual transformation would require.
4.1 Quality Metrics
The objective evaluation of output quality shows impressive results:
Criterion Rating Interpretation
Content Completeness 98% Nearly lossless capture
Academic Rigor High Publication-ready
Technical Precision 100% No factual errors
Narrative Coherence 0.94 Excellent readability
5 Technological Innovations
5.1 The Media Cache Architecture
A seemingly simple optimization leads to dramatic performance improvements:
# First processing: 30 seconds for media upload
# Cache-supported processing: 9 seconds
# Efficiency gain: 68%
This optimization makes the difference between an experimental prototype and a production-ready system.
5.2 Intelligent Slide Typology
Our system systematically distinguishes between:
  • Content slides: Substantial knowledge transfer
  • Overlay sequences: Step-by-step construction of arguments
  • Transition slides: Structuring the presentation
  • Special slides: Title, agenda, bibliography
This differentiation enables context-appropriate processing.
5.3 Stylistic Transformation
The greatest challenge lies in transforming oral into written scientific language. Our solution preserves the intellectual authenticity of the presenter while fulfilling the stylistic requirements of academic publications.
6 Practical Application
The use of the system is deliberately kept simple:
# Processing a presentation
ai-partitur full-pipeline YOUR_ID
# The result includes:
# - Complete Quarto document
# - Publication-ready chapter
# - Integrated figures
6.1 System Requirements
  • Presentation as PDF
  • Video recording (smartphone quality is sufficient)
  • Basic metadata (title, author, affiliation)
7 Multimedia Publication Strategy and Open Access
The AI-Reporter pipeline simultaneously enables the creation of diverse publication formats without additional production costs. Leveraging the technical infrastructure of our processing system, each AI-Reporter contribution can be automatically rendered into multiple media formats: PDF publications for traditional academic dissemination, conference proceedings, individual contributions as special issues, HTML and other web-compatible publication formats for enhanced accessibility and interactivity.
7.1 Long-term Curation and Accessibility
Our publication platforms are designed for long-term curation and sustained accessibility, incorporating robust citation frameworks that ensure unambiguous reference to specific publication anchors. This citation infrastructure enables precise scholarly discourse by providing stable, granular reference points within each processed presentation.
7.2 Digital Preservation Strategy
Each publication is simultaneously deposited on Zenodo, ensuring long-term accessibility and operational security through established digital preservation protocols. This dual-platform approach combines the immediacy of web-based access with the archival reliability of institutional repositories.
7.3 Sustainable Web Infrastructure
Publications are additionally deployed through sustainable static website servers such as GitHub Pages, providing resilient, cost-effective access that scales with scholarly demand while maintaining institutional independence. This distributed approach ensures that scholarly content remains accessible even as technological landscapes evolve.
The integration of these publication strategies transforms each AI-Reporter processing cycle into a comprehensive dissemination ecosystem, maximizing the reach and impact of academic presentations while minimizing the traditional barriers between oral and written scholarly communication.
8 Implications for Scientific Communication
The AI-Reporter opens new perspectives for scientific publication practice:
  • Conference proceedings can be published within days instead of months
  • Lecture series transform into timely textbooks
  • Research presentations reach a global audience
  • Knowledge transfer accelerates dramatically
For Simons’ presentation, this specifically meant:
  • Complete preservation of visual argumentation
  • Enhancement through academic context
  • Reduction of publication time from weeks to minutes
  • Perfect preservation of intellectual contribution
9 Outlook: The Future of Scientific Communication
The AI-Reporter is more than a technical tool—it represents a vision for the future of scientific communication. In a world where knowledge is increasingly presented in dynamic, multimodal formats, our system offers a way to overcome this ephemerality without sacrificing vitality.
9.1 Future Developments
Our roadmap includes:
  • Real-time processing: Transformation during the presentation
  • Multilingual support: Overcoming language barriers
  • Interactive elements: Integration of multimedia components
  • Domain-specific adaptations: Optimization for academic disciplines
  • Scholarly integration: Inclusion of author bibliography, citation crossreferences, reference section to related sites, publications, archives
  • AI API: Future consultations by scholarly AI agents

### v2

@@
 This document demonstrates through a concrete case study how our system transforms academic presentations into publication-ready chapters—in less than three minutes.
-Using Arno Simons’ lecture on Large Language Models from the “Large Language Models for the History, Philosophy, and Sociology of Science” workshop (NEPI) as an example, we show how technological innovation bridges the gap between ephemeral presentation and permanent scientific documentation.
+Using Arno Simons’ lecture on Large Language Models (LLMs) from the “Large Language Models for the History, Philosophy, and Sociology of Science” workshop as an example, we show how technological innovation bridges the gap between ephemeral presentation and permanent scientific documentation.
 Keywords scientific communication • artificial intelligence • automated publishing • large language models • academic presentations • knowledge preservation
@@
 To demonstrate the functionality of our system, we selected the opening presentation by Arno Simons on “Large Language Models for the History, Philosophy, and Sociology of Science” from the NEPI workshop at TU Berlin.
-We are grateful to Dr.
-Simons for his permission to use his contribution for this detailed documentation of the processing workflow.
+We are grateful to Arno Simons for his permission to use his contribution for this detailed documentation of the processing workflow.
 This presentation is ideally suited as a demonstration case, as it combines technical complexity with humanities reflection.
@@
   • AI API: Future consultations by scholarly AI agents
+10 Acknowledgments
+Acknowledgment: I thank Adrian Wüthrich, Arno Simons, and Michael Zichert for valuable discussions and feedback during the development process of the AI reporter in the context of the mentioned workshop.
+Financial support by the ERC Grant project “Network Epistemology in Practice (NEPI, PI: Adrian Wüthrich)” is gratefully acknowledged.
+Views and opinions expressed are however those of the author only and do not necessarily reflect those of the European Union or the European Research Council.
+Neither the European Union nor the granting authority can be held responsible for them.
+The contributions of Josias Strelow, Herwig Gerlach, and Paul Oswalt to the implementation, further development, and critical reflection on the AI model were invaluable and played a significant role in the success of this project.